OpenStudy (anonymous):

Write the equation of the horizontal line passing through the point (-1, -8).

OpenStudy (anonymous):

Pls help

OpenStudy (anonymous):

First, all horizontal lines have a slope of zero. If you use the point-slope formula: y-y \[y−y_{1}=m(x−x_{1})\]we will fill in the values accordingly. y−(−8)=0(x−(−1)) y+8=0 When you subtract 8 from both sides, you are left with y= -8. So, the equation of the horizontal line that passes through the point (-1, -8) is y= -8.

OpenStudy (anonymous):

Please disregard the "y-y" at the top.